# SURE-VQA: Systematic Understanding of Robustness Evaluation in Medical VQA Tasks

### v1

@@ C.4 Detailed Results of the Robustness Study
No Finetune 0.51 0.49 0.97 2.36 2.31 0.98 0.5 0.49 0.98 2.37 2.34 0.99 0.51 0.47 0.92 2.36 2.36 1
Full Finetune 0.75 0.75 1.01 3.41 3.44 1.01 0.72 0.77 1.07 3.27 3.51 1.07 0.71 0.79 1.12 3.3 3.45 1.05
Prompt 0.52$\pm$0.06 0.54$\pm$0.05 1.04$\pm$0.03 3.25$\pm$0.05 3.3$\pm$0.05 1.01$\pm$0.0 0.54$\pm$0.14 0.64$\pm$0.14 1.19$\pm$0.05 3.14$\pm$0.11 3.37$\pm$0.28 1.07$\pm$0.05 0.51$\pm$0.14 0.66$\pm$0.07 1.37$\pm$0.35 3.19$\pm$0.03 3.35$\pm$0.08 1.05$\pm$0.01
LoRA 0.75$\pm$0.01 0.76$\pm$0.0 1.02$\pm$0.01 3.35$\pm$0.11 3.41$\pm$0.06 1.02$\pm$0.01 0.71$\pm$0.03 0.79$\pm$0.02 1.12$\pm$0.02 3.32$\pm$0.04 3.58$\pm$0.01 1.08$\pm$0.01 0.73$\pm$0.01 0.78$\pm$0.01 1.07$\pm$0.0 3.36$\pm$0.04 3.54$\pm$0.1 1.05$\pm$0.02
$\mathtt{(IA)^{3}}$ 0.63$\pm$0.1 0.65$\pm$0.08 1.04$\pm$0.04 3.32$\pm$0.04 3.36$\pm$0.05 1.01$\pm$0.01 0.6$\pm$0.06 0.7$\pm$0.05 1.16$\pm$0.02 3.26$\pm$0.04 3.51$\pm$0.0 1.08$\pm$0.02 0.52$\pm$0.02 0.72$\pm$0.06 1.39$\pm$0.05 3.18$\pm$0.09 3.34$\pm$0.21 1.05$\pm$0.04
@@
No Finetune 0.49 0.48 0.97 2.34 2.37 1.01 0.49 0.48 0.97 2.37 2.39 1.01 0.51 0.46 0.89 2.33 2.49 1.07
Prompt 0.53$\pm$0.0 0.5$\pm$0.0 0.95$\pm$0.0 2.71$\pm$0.01 2.63$\pm$0.02 0.97$\pm$0.01 0.54$\pm$0.0 0.42$\pm$0.0 0.78$\pm$0.0 2.71$\pm$0.04 2.43$\pm$0.02 0.9$\pm$0.01 0.6$\pm$0.0 0.34$\pm$0.01 0.57$\pm$0.01 2.89$\pm$0.03 2.3$\pm$0.02 0.8$\pm$0.0
LoRA 0.53$\pm$0.0 0.5$\pm$0.0 0.95$\pm$0.0 2.74$\pm$0.01 2.64$\pm$0.01 0.97$\pm$0.0 0.54$\pm$0.0 0.41$\pm$0.0 0.76$\pm$0.0 2.75$\pm$0.02 2.43$\pm$0.02 0.88$\pm$0.01 0.59$\pm$0.01 0.35$\pm$0.02 0.6$\pm$0.04 2.93$\pm$0.04 2.25$\pm$0.02 0.77$\pm$0.01
$\mathtt{(IA)^{3}}$ 0.53$\pm$0.01 0.51$\pm$0.01 0.95$\pm$0.01 2.72$\pm$0.03 2.64$\pm$0.02 0.97$\pm$0.01 0.54$\pm$0.0 0.42$\pm$0.0 0.77$\pm$0.0 2.71$\pm$0.02 2.46$\pm$0.02 0.91$\pm$0.01 0.6$\pm$0.0 0.34$\pm$0.0 0.57$\pm$0.01 2.91$\pm$0.02 2.29$\pm$0.04 0.79$\pm$0.01
@@
Figure 12: Standard deviation between shifts vs. standard deviation between PEFT methods, not including full FT. The type of shift has a higher impact on the robustness than the PEFT method.
@@ Appendix D Corruption Study
We compared the realistic shifts defined for our datasets (R1) with artificial shifts, meaning image corruptions, to assess whether artificial shifts correspond to real-world shifts. The artificial data shifts were generated through image corruptions, including blur, Gaussian noise, and brightness adjustments. They were applied in different strengths (low, medium, and high). We used OpenCV for the image corruptions with the settings shown in Table 20.
@@
For this sample study, we used the LLaVA-Med model fine-tuned on the SLAKE dataset with the $\mathtt{(IA)^{3}}$method. The i.i.d. and OoD samples for realistic shifts were as previously described (R1). For artificial shifts, the i.i.d. train and test samples were identical to those used for realistic shifts, while OoD test samples were created by corrupting the i.i.d. test images with varying strengths of blur, brightness, and noise. Each corruption method was applied with a probability of $0.5$, with at least one corruption always being applied.
Table 21 shows the relative robustness results for both artificial and realistic shifts. The results show that both modality shift and question type shift exhibit lower relative robustness compared to all artificial shifts at low, medium, and high strengths. This suggests that artificial shifts, such as image corruption, fail to accurately represent the challenges posed by real-world, realistic shifts. The most prominent example here is the relative robustness of closed-ended questions under the question type shift (realistic shift), which is up to 96% compared to the realistic shift which only has 61%. The only exception where the realistic shift shows higher robustness is the question type shift on the open-ended questions, which is already nearly 100% on the realistic shift.
@@
Low Corruption 0.85$\pm$0.01 0.83$\pm$0.01 0.98$\pm$0.0 4.35$\pm$0.02 4.21$\pm$0.05 0.97$\pm$0.02 0.87$\pm$0.02 0.84$\pm$0.0 0.96$\pm$0.02 4.23$\pm$0.01 4.16$\pm$0.03 0.98$\pm$0.01
Medium Corruption 0.85$\pm$0.01 0.79$\pm$0.02 0.94$\pm$0.02 4.35$\pm$0.02 4.0$\pm$0.09 0.92$\pm$0.02 0.87$\pm$0.02 0.82$\pm$0.01 0.94$\pm$0.01 4.23$\pm$0.01 3.96$\pm$0.03 0.94$\pm$0.01
High Corruption 0.85$\pm$0.01 0.74$\pm$0.01 0.87$\pm$0.01 4.35$\pm$0.02 3.79$\pm$0.09 0.87$\pm$0.02 0.87$\pm$0.02 0.76$\pm$0.02 0.87$\pm$0.03 4.23$\pm$0.01 3.87$\pm$0.03 0.91$\pm$0.01
@@
Realistic Shift 0.85$\pm$0.01 0.64$\pm$0.07 0.75$\pm$0.08 4.35$\pm$0.02 3.4$\pm$0.15 0.78$\pm$0.04 0.87$\pm$0.02 0.53$\pm$0.06 0.61$\pm$0.06 4.23$\pm$0.01 4.21$\pm$0.27 0.99$\pm$0.07
Appendix E Multimodal Shifts
We conducted an ablation study on the OVQA dataset to evaluate the impact of a multimodal shift compared to the previously introduced unimodal shifts. This multimodal shift combines the Manifestation (Body Part) and Question Type Shifts reported in our experiments. Specifically, we defined the OoD set as samples featuring body part ”Leg” and question type ”Organ System”, with all other samples classified as i.i.d. As shown in Figure 13, the multimodal shift demonstrates the lowest robustness compared to unimodal shifts, which is expected given that multimodal shifts represent a more extreme divergence than their unimodal components.
Figure 13: Performance results on OVQA dataset with image organ shift, question type shift and multimodal shift which combines image organ shift and question type shift.

### v2

@@ C.5 Detailed Results of the Robustness Study
No Finetuning 0.51 0.5 0.97 1.81 1.77 0.97 0.51 0.5 0.99 1.86 1.77 0.95 0.52 0.48 0.93 1.86 1.8 0.97
Full Finetuning 0.75 0.75 1 3.41 3.43 1.01 0.73 0.78 1.07 3.26 3.54 1.09 0.71 0.79 1.1 3.27 3.51 1.07
Prompt Tuning 0.48 +/- 0.09 0.48 +/- 0.08 1.01 +/- 0.04 3.27 +/- 0.04 3.28 +/- 0.05 1.0 +/- 0.0 0.53 +/- 0.15 0.63 +/- 0.16 1.2 +/- 0.06 3.13 +/- 0.11 3.37 +/- 0.28 1.07 +/- 0.05 0.49 +/- 0.17 0.62 +/- 0.13 1.33 +/- 0.29 3.19 +/- 0.03 3.33 +/- 0.08 1.05 +/- 0.01
LoRA 0.75 +/- 0.0 0.76 +/- 0.0 1.02 +/- 0.01 3.38 +/- 0.07 3.41 +/- 0.04 1.01 +/- 0.01 0.71 +/- 0.03 0.79 +/- 0.02 1.11 +/- 0.02 3.31 +/- 0.03 3.6 +/- 0.01 1.09 +/- 0.01 0.73 +/- 0.01 0.79 +/- 0.01 1.07 +/- 0.0 3.34 +/- 0.03 3.58 +/- 0.06 1.07 +/- 0.01
IA3 0.61 +/- 0.15 0.63 +/- 0.12 1.04 +/- 0.06 3.33 +/- 0.02 3.36 +/- 0.04 1.01 +/- 0.01 0.61 +/- 0.05 0.7 +/- 0.05 1.15 +/- 0.02 3.25 +/- 0.03 3.52 +/- 0.03 1.08 +/- 0.02 0.58 +/- 0.1 0.74 +/- 0.06 1.3 +/- 0.15 3.2 +/- 0.09 3.41 +/- 0.2 1.06 +/- 0.03
(b) Results for the non-medical base model.
@@
No Finetuning 0.53 0.54 1.03 1.58 1.6 1.01 0.51 0.55 1.07 1.62 1.67 1.03 0.52 0.55 1.07 1.61 1.68 1.05
Full Finetuning 0.59 0.61 1.02 2.6 2.68 1.03 0.5 0.55 1.09 2.38 2.6 1.09 0.53 0.33 0.62 2.62 2.37 0.91
Prompt Tuning 0.66 +/- 0.02 0.66 +/- 0.03 1.01 +/- 0.02 3.14 +/- 0.06 3.15 +/- 0.04 1.0 +/- 0.01 0.55 +/- 0.05 0.63 +/- 0.02 1.16 +/- 0.1 2.98 +/- 0.05 3.18 +/- 0.23 1.07 +/- 0.06 0.63 +/- 0.02 0.51 +/- 0.15 0.8 +/- 0.21 3.1 +/- 0.01 3.05 +/- 0.09 0.99 +/- 0.03
LoRA 0.74 +/- 0.01 0.75 +/- 0.01 1.02 +/- 0.01 3.43 +/- 0.0 3.47 +/- 0.03 1.01 +/- 0.01 0.69 +/- 0.02 0.77 +/- 0.01 1.12 +/- 0.02 3.37 +/- 0.02 3.68 +/- 0.04 1.09 +/- 0.01 0.69 +/- 0.06 0.77 +/- 0.05 1.12 +/- 0.03 3.37 +/- 0.03 3.6 +/- 0.08 1.07 +/- 0.02
IA3 0.69 +/- 0.06 0.72 +/- 0.05 1.04 +/- 0.02 3.31 +/- 0.02 3.35 +/- 0.01 1.01 +/- 0.01 0.68 +/- 0.03 0.78 +/- 0.01 1.14 +/- 0.03 3.21 +/- 0.02 3.51 +/- 0.03 1.09 +/- 0.0 0.65 +/- 0.09 0.71 +/- 0.12 1.1 +/- 0.04 3.15 +/- 0.16 3.33 +/- 0.08 1.06 +/- 0.05
(a) Results for the medical base model.
Gender Shift Ethnicity Shift Age Shift
OoD: Female OoD: Non-white OoD: Young
No Finetuning 0.51 0.49 0.96 1.5 1.48 0.99 0.51 0.5 0.98 1.56 1.49 0.96 0.53 0.47 0.89 1.58 1.48 0.94
Prompt Tuning 0.54 +/- 0.0 0.52 +/- 0.0 0.95 +/- 0.0 2.74 +/- 0.01 2.63 +/- 0.01 0.96 +/- 0.01 0.55 +/- 0.01 0.42 +/- 0.0 0.77 +/- 0.0 2.73 +/- 0.03 2.46 +/- 0.01 0.9 +/- 0.01 0.6 +/- 0.0 0.34 +/- 0.01 0.57 +/- 0.01 2.9 +/- 0.02 2.31 +/- 0.02 0.79 +/- 0.01
LoRA 0.55 +/- 0.0 0.51 +/- 0.0 0.94 +/- 0.0 2.75 +/- 0.01 2.64 +/- 0.01 0.96 +/- 0.0 0.55 +/- 0.0 0.42 +/- 0.0 0.76 +/- 0.0 2.74 +/- 0.03 2.45 +/- 0.02 0.89 +/- 0.01 0.59 +/- 0.01 0.35 +/- 0.02 0.6 +/- 0.04 2.9 +/- 0.05 2.25 +/- 0.01 0.78 +/- 0.01
IA3 0.55 +/- 0.01 0.52 +/- 0.01 0.95 +/- 0.01 2.75 +/- 0.02 2.65 +/- 0.01 0.96 +/- 0.01 0.55 +/- 0.0 0.42 +/- 0.0 0.77 +/- 0.01 2.72 +/- 0.02 2.48 +/- 0.02 0.91 +/- 0.01 0.6 +/- 0.0 0.34 +/- 0.01 0.56 +/- 0.01 2.92 +/- 0.01 2.32 +/- 0.03 0.8 +/- 0.01
(b) Results for the non-medical base model.
Gender Shift Ethnicity Shift Age Shift
OoD: Female OoD: Non-white OoD: Young
No Finetuning 0.43 0.42 0.98 1.21 1.22 1.01 0.43 0.38 0.9 1.27 1.27 1 0.45 0.37 0.82 1.27 1.29 1.02
Prompt Tuning 0.54 +/- 0.0 0.51 +/- 0.0 0.95 +/- 0.0 2.72 +/- 0.01 2.64 +/- 0.01 0.97 +/- 0.0 0.53 +/- 0.03 0.45 +/- 0.05 0.85 +/- 0.14 2.72 +/- 0.0 2.46 +/- 0.04 0.9 +/- 0.02 0.5 +/- 0.08 0.49 +/- 0.12 1.03 +/- 0.44 2.9 +/- 0.02 2.32 +/- 0.02 0.8 +/- 0.0
LoRA 0.53 +/- 0.01 0.51 +/- 0.01 0.96 +/- 0.03 2.7 +/- 0.01 2.63 +/- 0.02 0.97 +/- 0.01 0.54 +/- 0.0 0.41 +/- 0.0 0.76 +/- 0.01 2.72 +/- 0.02 2.5 +/- 0.01 0.92 +/- 0.01 0.6 +/- 0.0 0.34 +/- 0.0 0.57 +/- 0.01 2.85 +/- 0.02 2.29 +/- 0.06 0.8 +/- 0.01
IA3 0.46 +/- 0.11 0.46 +/- 0.12 1.01 +/- 0.02 2.72 +/- 0.02 2.67 +/- 0.01 0.98 +/- 0.01 0.54 +/- 0.0 0.45 +/- 0.04 0.84 +/- 0.07 2.71 +/- 0.0 2.5 +/- 0.0 0.92 +/- 0.0 0.59 +/- 0.01 0.34 +/- 0.02 0.58 +/- 0.04 2.87 +/- 0.01 2.32 +/- 0.02 0.81 +/- 0.01
@@
Figure 18: Standard deviation between shifts vs. standard deviation between PEFT methods for the medical base model, not including full FT. The type of shift has a higher impact on the robustness than the PEFT method.
C.5.1 Statistical analysis
In this section, we confirm two of the main findings in our robustness study by a detailed statistical analysis.
No FT method consistently outperforms others in terms of robustness.
To prove this in addition to the rank analysis in Figure 8b), we run a pairwise Welch’s t-test between the FT methods across the different shifts and correct for multiple testing. Each test-set is bootstrapped such that we have $100$ robustness measures per experiment. We then report the Win/Loss Matrix with a p value of $p=0.05$ (Bahri et al. (2022)) to see which FT method significantly outperforms another FT method significantly on a shift. The results are shown in Figure 19. While there is a slight trend indicating that Full FT is less likely to outperform other methods and is more frequently outperformed by them, the remaining methods show very similar performance levels. Overall, this analysis suggests that there is no clear winner of a method that more often outperforms others and is at the same time not often outperformed.
Figure 19: Win/Loss Matrix for a pairwise comparison between FT methods. The entries show cases where one FT method outperforms the other significantly, measures by a Welch’s t-test with $p=0.05$. A good method would achieve high values in its row (often outperforming others) and low values in its column (not often outperformed by others).
Robustness trends are more stable across FT methods than across distribution shifts.
For this comparison, we conduct a one-way ANOVA on the robustness scores to compare the variability across FT methods versus across distribution shift types. The corresponding F-values and p-values are shown in Table 20. Confirming Figure 8c), we see that generally, the variance between the shifts is significantly higher than between methods ($F>1,p<0.05$). The only outlier is OVQA with a one shift being instable for Full FT, but this trend changes when excluding Full FT as also seen in Figure 18.
Table 20: One way ANOVA to analyze the variance between shifts and variance between methods.
[…]
F-value p-value F-Value p-value
SLAKE $19.16$ $7.17\times 10^{-5}$ $16.83$ $0.0008$
OVQA $1.11$ $0.38$ $21.71$ $0.0003$
MIMIC $5.92$ $0.002$ $5.84$ $0.006$
@@ Appendix D Corruption Study
This section discusses the details of the corruption study showing the empirical confirmation of R1 (see section 3.3.1). Figure 20 shows the visualization of the corruption study pipeline, illustrating how artificial and realistic OoD test sets are generated and used to evaluate model robustness. We use OpenCV for image corruptions with the settings shown in Table 21.
@@
Table 22 shows the relative robustness results for both artificial and realistic shifts. The results show that both modality shift and question type shift exhibit lower relative robustness compared to all artificial shifts at low, medium, and high strengths. This suggests that artificial shifts, such as image corruption, fail to accurately represent the challenges posed by real-world, realistic shifts. The most prominent example here is the relative robustness of closed-ended questions under the artificial corruption shift which is up to 96% compared to the realistic shift (question type) which only has 61%. The only exception where the realistic shift shows higher robustness is the question type shift on the open-ended questions.
@@
Low Corruption 0.85 $\pm$ 0.01 0.83 $\pm$ 0.01 0.98 $\pm$ 0.0 4.36 $\pm$ 0.03 4.25 $\pm$ 0.04 0.98 $\pm$ 0.02 0.87 $\pm$ 0.02 0.84 $\pm$ 0.0 0.96 $\pm$ 0.02 4.27 $\pm$ 0.01 4.2 $\pm$ 0.03 0.98 $\pm$ 0.01
Medium Corruption 0.85 $\pm$ 0.01 0.79 $\pm$ 0.02 0.94 $\pm$ 0.02 4.36 $\pm$ 0.03 4.06 $\pm$ 0.07 0.93 $\pm$ 0.02 0.87 $\pm$ 0.02 0.82 $\pm$ 0.01 0.94 $\pm$ 0.01 4.27 $\pm$ 0.01 4.02 $\pm$ 0.04 0.94 $\pm$ 0.01
High Corruption 0.85 $\pm$ 0.01 0.74 $\pm$ 0.01 0.87 $\pm$ 0.01 4.36 $\pm$ 0.03 3.84 $\pm$ 0.08 0.88 $\pm$ 0.03 0.87 $\pm$ 0.02 0.76 $\pm$ 0.02 0.87 $\pm$ 0.03 4.27 $\pm$ 0.01 3.94 $\pm$ 0.03 0.92 $\pm$ 0.01
@@
Realistic Shift 0.85 $\pm$ 0.01 0.64 $\pm$ 0.07 0.75 $\pm$ 0.08 4.36 $\pm$ 0.03 3.46 $\pm$ 0.12 0.79 $\pm$ 0.03 0.87 $\pm$ 0.02 0.53 $\pm$ 0.06 0.61 $\pm$ 0.06 4.27 $\pm$ 0.01 4.15 $\pm$ 0.22 0.97 $\pm$ 0.05
Figure 20: Overview of the Corruption Study Pipeline. The left side of the figure illustrates the training phase, where the model is fine-tuned on an i.i.d. training set. On the right, the inference phase depicts how two types of OoD test sets (artificial and realistic) are generated and used to evaluate robustness. For both the artificial and realistic shifts, the same i.i.d. test set is used. The corrupted OoD test set is generated by applying described artificial corruptions to the i.i.d. test images. The realistically shifted OoD test set is constructed using the data shifts introduced in section 3.1. Robustness is then measured by comparing the model’s performance on the i.i.d. test set to its performance on each respective OoD set.
Appendix E Swapped Shifts
We conduct an ablation study on the SLAKE dataset to see how changing the i.i.d. and OoD set influences the robustness. Thereby, we use the same type of shifts as described in section 3.1 (Modality and Question Type), but with different i.i.d. and OoD sets. The concrete realization of the swapped shifts are shown in Table 23.
Table 23: Original and Swapped Shifts on the SLAKE Dataset.
Shift Type Original Shift Swapped Shift
i.i.d. OoD i.i.d. OoD
Modality CT/MRI X-Ray CT/X-Ray MRI
Question Type All questions except Size Size All questions except Position Position
Figure 21: Results on the SLAKE dataset with swapped shifts, meaning the i.i.d. and OoD splits are changed. For each shift (Original Shifts, Swapped Shifts) the top row shows the i.i.d. and OoD performance on Open- and Closed-Ended questions, lower left shows the Relative Robustness, and lower right shows the Inter-Shift vs. Inter-Method variability.
The results are shown in Figure 21. Overall, there is a swap in closed-ended and open-ended robustness: While for the original shifts, the models showed less robustness on the closed-ended questions, in the swapped shifts, they show less robustness on the open-ended questions. As also noted in section 3.4, this arises from the related question patterns between certain OoD questions and the i.i.d. training data. Although the OoD questions are not explicitly seen during training, similar questions in the training set can provide enough contextual information for effective generalization. In this case, performance on the closed-ended OoD questions is higher because training examples like "Is this a study of the [body region]?" implicitly support OoD questions such as "Do the organs in the image exist in the [body region]?" This observation confirms one of our main findings: the type of shift has a greater impact on robustness than the choice of fine-tuning method. To achieve more reliable and generalizable performance, it is therefore more effective to focus on increasing the diversity of training data—so that it better reflects the target population—rather than solely optimizing the fine-tuning strategy.
Appendix F Multimodal Shifts
We conduct an ablation study on the OVQA dataset to evaluate the impact of a multimodal shift compared to the previously introduced unimodal shifts. Similar to the corruption study (Appendix D), this ablation study is performed on the medical base model with $\mathtt{(IA)^{3}}$as PEFT method. This multimodal shift combines the manifestation (Body Part) and question type shifts reported in our experiments. Specifically, we define the OoD set as samples featuring body part "Leg" and question type "Organ System", with all other samples classified as i.i.d. As shown in Figure 22, the multimodal shift demonstrates the lowest robustness compared to unimodal shifts, which is expected given that multimodal shifts represent a more extreme divergence than the unimodal ones.
Figure 22: Performance results on OVQA dataset with image organ shift, question type shift, and multimodal shift which combines image organ shift and question type shift.